# ScribeAgent: Towards Specialized Web Agents Using Production-Scale Workflow Data

### v1

@@ Abstract
Large Language Model (LLM) agents are rapidly improving to handle increasingly complex web-based tasks. Most of these agents rely on general-purpose, proprietary models like GPT-4 and focus on designing better prompts to improve their planning abilities. However, general-purpose LLMs are not specifically trained to understand specialized web contexts such as HTML, and they often struggle with long-horizon planning. We explore an alternative approach that fine-tunes open-source LLMs using production-scale workflow data collected from over 250 domains corresponding to 6 billion tokens. This simple yet effective approach shows substantial gains over prompting-based agents on existing benchmarks—ScribeAgent achieves state-of-the-art direct generation performance on Mind2Web and improves the task success rate by 14.1% over the previous best text-only web agents on WebArena. We further perform detailed ablation studies on various fine-tuning design choices and provide insights into LLM selection, training recipes, context window optimization, and effect of dataset sizes.
@@ 1 Introduction
Large language model (LLM) agents have advanced significantly in web navigation. They can carry out user-specified tasks in multiple steps by reasoning on their own what actions to take and what external resources to interface with. Recent studies (Zheng et al., 2024; Lai et al., 2024; Zhang et al., 2024) have shown that, with better planning and exploration strategies, LLM agents can independently solve various web tasks ranging from simple navigation, such as locating a specific Wikipedia page, to more complex operations, such as booking flights or restaurants.
@@
With access to this production-scale dataset, we develop ScribeAgent, the first family of specialized, single-stage LLM agents capable of directly generating the next step based on the website’s DOM and action history. This is in contrast with previous fine-tuned agents that require multiple stages to produce an action, e.g., first narrowing down to a set of target element candidates and then selecting one from the candidates (Deng et al., 2023). To evaluate the capacity and generalization ability of ScribeAgent, we test it on public benchmarks without any further training. ScribeAgent significantly outperforms existing GPT-4-based and multi-stage agents. Notably, the 32B-parameter ScribeAgent-Large achieves state-of-the-art direct generation performance on Mind2Web (Deng et al., 2023), with step success rate surpassing the baselines by 5-10% across all test sets. On the end-to-end task execution benchmark WebArena (Zhou et al., 2024), our 7B ScribeAgent-Small improves the previous best task success rate from 37.2% to 51.3%, marking the highest performance among text-only LLM agents.
@@ 2 Related Work
Prompting-based agent frameworks. The majority of web agent works reuse existing LLMs and propose different prompting strategies to improve action prediction. One line of research focuses on exploiting previous experience via self-feedback (Sun et al., 2023) or in-context demonstrations (Fu et al., 2024; Zheng et al., 2024; Wang et al., 2024b; Ou et al., 2024; Shen et al., 2024b). A separate line of work centers around encouraging exploration by including external evaluators (Pan et al., 2024), using synthesized instructions (Murty et al., 2024), or applying more sophisticated search algorithms like stack (Sodhi et al., 2024), best-first tree search (Koh et al., 2024), or Monte Carlo Tree Search (Zhang et al., 2024). Despite the research efforts, these prompting methods rely heavily on the quality of the LLM used. Open-source models such as LLaMA (Dubey et al., 2024), Code LLaMA (Rozière et al., 2024), and Flan-T5 (Chung et al., 2022) generally underperform proprietary models like GPT-4. However, fine-tuning proprietary LLMs can often be costly and challenging, as it is restricted to being done through APIs. This implies an opportunity for enhancing open-source LLMs to match or outperform proprietary agents.
Fine-tuning-based web agents. Compared to developing better reasoning and planning frameworks, comparatively less attention has been given to optimizing the LLMs themselves to better handle web environments. Due to the difficulty of directly generating a single target element from the raw HTML, which often contains thousands of elements, existing work mostly focuses on multi-stage prediction. MindAct (Deng et al., 2023) proposes a two-stage pipeline that first uses a small LM to filter the web elements and then uses a more powerful LM to select from the filtered elements in a multi-choice question answering format. Both LMs can be fine-tuned using the Mind2Web dataset. WebAgent (Gur et al., 2023) uses HTML-5 to first process the HTML and then fine-tunes a 540B Flan-UPalm to generate code for controlling web pages. More recently, AutoWebGLM (Lai et al., 2024) trains a single ChatGLM3 6B (GLM et al., 2024) using a combination of curriculum learning, reinforcement learning, and rejection sampling fine-tuning. Despite the complicated training and inference procedures, these methods often underperform agents that prompt GPT-4. In contrast, our work shows that given sufficient high-quality workflow data, fine-tuning a single LLM can achieve strong performance. We note that the newly released OpenAI o1 (OpenAI, 2024c) can be viewed as a specialized agent with a complicated planning framework. Nonetheless, we show in Section 4.1 that ScribeAgent outperforms o1-preview by a large margin on our proprietary dataset. Moreover, while none of the training details for o1 have been released, our work provides valuable insights into data preprocessing and fine-tuning.
@@ 3.2.4 Exploring the Design Space
Pretrained LLM Selection. Intuitively, the quality of a fine-tuned web agent should be relevant to the quality of the pretained LLM. We identify two axes that are crucial to performance—model architecture and model size—and explore seven open-source LLMs spanning these axes: Llama 3.1 8B (Dubey et al., 2024), Mistral 7B (MistralAI, 2023), Mixtral 8x7B (MistralAI, 2024b), Qwen2 7B (Yang et al., 2024), Qwen2 57B (Yang et al., 2024), Qwen2.5 14B (Team, 2024), Qwen2.5 32B (Team, 2024), and Codestral 22B (MistralAI, 2024a). We fine-tune these models with 1 billion training tokens and evaluate their performance on the test split of the dataset we collected.
@@ 4.3 End-to-End Task Execution on WebArena
We apply the above pipeline to solve the WebArena tasks. In stage 3, GPT-4o maps the agent’s output in HTML to the accessibility tree format, which is then processed by the WebArena environment. To further improve performance, we allow ScribeAgent to generate multiple actions in stage 2 and select the one with the highest confidence using majority vote and GPT-4o analysis. Given the large number of tasks in WebArena and the evaluation costs associated with using a multi-agent system, we limited our testing to ScribeAgent-Small. More details about evaluating ScribeAgent on WebArena can be found in Appendix A.5.
Table 7: Task success rates (SR) on WebArena and score breakdown on five web domains. ScribeAgent consistently outperforms all considered baselines, often improving the previous-best results by more than 10%. Note that we only test ScribeAgent-Small due to the large number of tasks in WebArena and the evaluation costs associated with the multi-agent system.
@@
Multi-Agent System (Ours) ScribeAgent-Small + GPT4o 51.3 48.1 35.5 70.2 58.8 51.9
@@
The results are shown in Table 7. Compared with existing text-only baselines, ScribeAgent-Small augmented with GPT-4o obtains the highest task success rate in all five categories, leading to 14.1% performance improvements in total success rate over the previous-best WebPilot results. In particular, on Reddit and GitLab tasks where the domains are more realistic and thus closer to the ones in our training data, ScribeAgent demonstrates stronger generalization ability and higher task success rates than in other domains. Despite known issues with combobox selection and the absence of scroll actions in our training data, our agent effectively navigates these challenges through strategic keyboard actions. Further details are provided in Appendix A.5.2.
To better understand the contribution of ScribeAgent to the multi-agent system, we perform an ablation study that leverages GPT-4o for all four-stages of the proposed pipeline. As shown in Table 8, using ScribeAgent consistently outperforms only using GPT-4o, and the GPT-4o-only setting is less effective than existing agents like WebPilot. This shows that our strong performance on WebArena can be mostly attributed to the action generation process of ScribeAgent. Apart from getting better results, the multi-agent system is cheaper than using GPT-4o alone, as calling ScribeAgent to generate a next action incurs negligible cost as it is served locally.
@@
Multi-Agent ScribeAgent-Small + GPT4o 51.3 48.1 35.5 70.2 58.8 51.9
Table 9: Task success rates on a subset of WebArena. The numbers after the domains indicate the number of tasks considered. All model are used along with GPT-4o to formulate the multi-agent system. We see that the general trends agree with what we found on our proprietary dataset.

### v2

@@ Abstract
Large Language Model (LLM) agents are rapidly improving to handle increasingly complex web-based tasks. Most of these agents rely on general-purpose, proprietary models like GPT-4 and focus on designing better prompts to improve their planning abilities. However, general-purpose LLMs are not specifically trained to understand specialized web contexts such as HTML, and they often struggle with long-horizon planning. We explore an alternative approach that fine-tunes open-source LLMs using production-scale workflow data collected from over 250 domains corresponding to 6 billion tokens. This simple yet effective approach shows substantial gains over prompting-based agents on existing benchmarks—ScribeAgent achieves state-of-the-art direct generation performance on Mind2Web and improves the task success rate by 7.3% over the previous best text-only web agents on WebArena. We further perform detailed ablation studies on various fine-tuning design choices and provide insights into LLM selection, training recipes, context window optimization, and effect of dataset sizes.
@@ 1 Introduction
Large language model (LLM) agents have advanced significantly in web navigation. They can carry out user-specified tasks in multiple steps by reasoning on their own what actions to take and what external resources to interface with. Recent studies (Zheng et al., 2024; Lai et al., 2024; Zhang et al., 2024; Song et al., 2024) have shown that, with better planning and exploration strategies, LLM agents can independently solve various web tasks ranging from simple navigation, such as locating a specific Wikipedia page, to more complex operations, such as booking flights or restaurants.
@@
With access to this production-scale dataset, we develop ScribeAgent, the first family of specialized, single-stage LLM agents capable of directly generating the next step based on the website’s DOM and action history. This is in contrast with previous fine-tuned agents that require multiple stages to produce an action, e.g., first narrowing down to a set of target element candidates and then selecting one from the candidates (Deng et al., 2023). To evaluate the capacity and generalization ability of ScribeAgent, we test it on public benchmarks without any further training. ScribeAgent significantly outperforms existing GPT-4-based and multi-stage agents. Notably, the 32B-parameter ScribeAgent-Large achieves state-of-the-art direct generation performance on Mind2Web (Deng et al., 2023), with step success rate surpassing the baselines by 5-10% across all test sets. On the end-to-end task execution benchmark WebArena (Zhou et al., 2024), our 7B ScribeAgent-Small improves the previous best task success rate from 45.7% to 51.3%. ScribeAgent-Large achieves an even better task success rate of 53%, marking the highest performance among text-only LLM agents.
@@ 2 Related Work
Prompting-based agent frameworks. The majority of web agent works reuse existing LLMs and propose different prompting strategies to improve action prediction. One line of research focuses on exploiting previous experience via self-feedback (Sun et al., 2023) or in-context demonstrations (Fu et al., 2024; Zheng et al., 2024; Wang et al., 2024b; Ou et al., 2024; Shen et al., 2024b). A separate line of work centers around encouraging exploration by including external evaluators (Pan et al., 2024), using synthesized instructions (Murty et al., 2024b), or applying more sophisticated search algorithms like stack (Sodhi et al., 2024), best-first tree search (Koh et al., 2024), or Monte Carlo Tree Search (Zhang et al., 2024). Despite the research efforts, these prompting methods rely heavily on the quality of the LLM used. Open-source models such as LLaMA (Dubey et al., 2024), Code LLaMA (Rozière et al., 2024), and Flan-T5 (Chung et al., 2022) generally underperform proprietary models like GPT-4. However, fine-tuning proprietary LLMs can often be costly and challenging, as it is restricted to being done through APIs. This implies an opportunity for enhancing open-source LLMs to match or outperform proprietary agents.
Fine-tuning-based web agents. Compared to developing better reasoning and planning frameworks, comparatively less attention has been given to optimizing the LLMs themselves to better handle web environments. Due to the difficulty of directly generating a single target element from the raw HTML, which often contains thousands of elements, existing work mostly focuses on multi-stage prediction. MindAct (Deng et al., 2023) proposes a two-stage pipeline that first uses a small LM to filter the web elements and then uses a more powerful LM to select from the filtered elements in a multi-choice question answering format. Both LMs can be fine-tuned using the Mind2Web dataset. WebAgent (Gur et al., 2023) uses HTML-5 to first process the HTML and then fine-tunes a 540B Flan-UPalm to generate code for controlling web pages. More recently, AutoWebGLM (Lai et al., 2024) trains a single ChatGLM3 6B (GLM et al., 2024) using a combination of curriculum learning, reinforcement learning, and rejection sampling fine-tuning. NNetnav (Murty et al., 2024a) leverages synthetic demonstrations collected by zero-shot LLM exploration and hindsight relabeling to fine-tune open-source models into web agents. Despite the complicated training and inference procedures, these methods often underperform agents that prompt GPT-4. In contrast, our work shows that given sufficient high-quality workflow data, fine-tuning a single LLM can achieve strong performance. We note that the newly released OpenAI o1 (OpenAI, 2024c) can be viewed as a specialized agent with a complicated planning framework. Nonetheless, we show in Section 4.1 that ScribeAgent outperforms o1-preview by a large margin on our proprietary dataset. Moreover, while none of the training details for o1 have been released, our work provides valuable insights into data preprocessing and fine-tuning.
@@ 3.2.4 Exploring the Design Space
Pretrained LLM Selection. Intuitively, the quality of a fine-tuned web agent should be relevant to the quality of the pretained LLM. We identify two axes that are crucial to performance—model architecture and model size—and explore seven open-source LLMs spanning these axes: Llama 3.1 8B (Dubey et al., 2024), Mistral 7B (MistralAI, 2023), Mixtral 8x7B (MistralAI, 2024b), Qwen2 7B (Yang et al., 2024a), Qwen2 57B (Yang et al., 2024a), Qwen2.5 14B (Team, 2024), Qwen2.5 32B (Team, 2024), and Codestral 22B (MistralAI, 2024a). We fine-tune these models with 1 billion training tokens and evaluate their performance on the test split of the dataset we collected.
@@ 4.3 End-to-End Task Execution on WebArena
We apply the above pipeline to solve the WebArena tasks. In stage 3, GPT-4o maps the agent’s output in HTML to the accessibility tree format, which is then processed by the WebArena environment. To further improve performance, we allow ScribeAgent to generate multiple actions in stage 2 and select the one with the highest confidence using majority vote and GPT-4o analysis. More details about evaluating ScribeAgent on WebArena can be found in Appendix A.5.
Table 7: Task success rates (SR) on WebArena and score breakdown on five web domains. ScribeAgent consistently outperforms considered text-only baselines, often improving the previous-best results by more than 5%. Note that we only test ScribeAgent-Small due to the large number of tasks in WebArena and the evaluation costs associated with the multi-agent system.
@@
Broswing+API Hybrid Agent GPT-4o 35.8 25.7 41.2 28.3 44.4 45.9
AgentOccam with Judge GPT-4-Turbo 45.7 43.3 46.2 67.0 38.9 52.3
Multi-Agent System (Ours) ScribeAgent-Small + GPT-4o 51.3 48.1 35.5 70.2 58.8 51.9
ScribeAgent-Large + GPT-4o 53.0 45.8 37.9 73.7 59.7 56.3
@@
The results are shown in Table 7. Compared with existing text-only baselines, ScribeAgent augmented with GPT-4o obtains the highest task success rate in 4 of 5 categories, leading to 7.3% performance improvements in total success rate over the previous-best GPT-4-Turbo-based AgentOccam (Yang et al., 2024b). In particular, on Reddit and GitLab tasks where the domains are more realistic and thus closer to the ones in our training data, ScribeAgent demonstrates stronger generalization ability and higher task success rates than in other domains. Despite known issues with combobox selection and the absence of scroll actions in our training data, our agent effectively navigates these challenges through strategic keyboard actions. More details are provided in Appendix A.5.2.
To better understand the contribution of ScribeAgent to the multi-agent system, we perform an ablation study that leverages GPT-4o for all four stages of the proposed pipeline. Given the large number of tasks in WebArena and the evaluation costs associated with using a multi-agent system, we perform our ablation studies using ScribeAgent-Small. As shown in Table 8, using ScribeAgent consistently outperforms only using GPT-4o, and the GPT-4o-only setting is less effective than existing agents like WebPilot. This shows that our strong performance on WebArena can be mainly attributed to the action generation process of ScribeAgent. Apart from getting better results, the multi-agent system is cheaper than using GPT-4o alone, as calling ScribeAgent to generate a next action incurs negligible cost as it is served locally.
@@
Multi-Agent ScribeAgent-Small + GPT-4o 51.3 48.1 35.5 70.2 58.8 51.9
Table 9: Task success rates on a subset of WebArena. The numbers after the domains indicate the number of tasks considered. All models are used along with GPT-4o to formulate the multi-agent system. We see that the general trends agree with what we found on our proprietary dataset.